OpenStudy (anonymous):

HELP!!!! Determine the equation of joint variation. Then solve for the missing value. x varies directly with y and z. x = 40 when y = 10 and z = 20. Find x when y = 30 and z = 40.

OpenStudy (anonymous):

x = kyz plugging in 40 = k*10*20 k = 40/200 = 1/5 equation : x = yz/5 plugging in x = 30*40/5 = 240 answer: 240

OpenStudy (anonymous):

Thank you very much!!!!